# MgFNO: Multi-grid Architecture Fourier Neural Operator for Parametric Partial Differential Equations

### v2

@@ 3.2 Compensation Methods of Low-Frequency Tendency
To address the issues caused by slow convergence at high frequencies and speed up the training of FNO, several methods have been proposed. Jagtap and Ameya D [27] used activation functions $\sigma(\mu ax)$ instead of $\sigma(\mu x)$. Here, $\mu$ is a fixed scaling factor, and $a$ is a trainable parameter for all neurons. Cai and Wei [28] introduced PhaseDNN, which converts the high-frequency components of data to the low-frequency spectrum for learning, and then converts the learned frequency components back to the original high-frequency spectrum. Subsequently, the multiscale DNN method (MscaleDNN) was proposed [22, 29], which can be understood as Fourier expansion only in the radial direction. The frequency space conversion can be achieved through scaling, which is equivalent to inverse scaling in spatial space. In this article, we utilize the conversion of coarse and fine grids and MscaleDNN to address the slow convergence issue of high-frequency errors in FNO. Next, let us review the ideas and structure of MscaleDNN in regular DNNs.
@@
A common MscaleDNN structure is shown in Figure 7. We use neural networks successively in physical space to learn the stretched sub-functions $f_{i}^{(scale)}(\alpha_{i}x)$. Furthermore, it is mentioned in [22] that when constructing MscaleDNN, activation functions with compact support should be selected. In MgFNO, we select the following function $\phi(x)$ as our activation function.
@@
Figure 7: : A common MscaleDNN structure. Reprinted from Liu et al. [22]
On the other hand, looking back at the MG method we mentioned in Section 2, it is worth noting that iterative methods (2.1) tend to eliminate high-frequency error components rather than low-frequency ones. Therefore, we may employ a straightforward approach to remove these low-frequency error components. As it is shown in figure 8, for the MG method on a coarser mesh, the low-frequency components transform into high-frequency ones. If there is an error on one side of the mesh and a different error on the other side of the mesh, once we coarsen the cells and then bring them next to each other, the variation between those two positions is now a high spatial frequency. Then we can remove high-frequency components with a few iterations of iterative methods. Due to the opposite convergence tendency of DNNs and iterative methods in the frequency domain, we have decided to apply a multi-grid architecture similar to the MG method to FNO, but this time we need to transfer the coarse grid to the fine grid. We hope that through the design of this multi-grid architecture, FNO can overcome the high-frequency issues and accelerate the model training process.
@@ 4.2 Architecture of MgFNO
Due to the FNO adhering to the F-Principle, we suggest that mitigating the high-frequency errors of FNO could further enhance its overall performance. To accomplish this, we propose MgFNO, a multigrid framework that trains multiple FNOs sequentially. There are various multigrid algorithms available, such as the V-cycle and W-cycle. For the convenience of model construction, we employ a three-layer V-cycle multigrid architecture to build the model, and the specific algorithm of MgFNO is depicted in Algorithm 2.
@@ 5 Numerical Implementation and Results
Data Set and Result. Firstly, we set the viscosity coefficient $\nu=0.1$. Then, for initial condition $u_{0}(x)$, we set $u_{0}\sim\mu$ where $\mu=\mathcal{N}(0,625(-\Delta+25I)^{-2})$. Finally, we solve the equation using a split step method where the heat equation part is solved exactly in Fourier space then the non-linear part is advanced, again in Fourier space, using a very fine forward Euler method. We solve on a spatial mesh with resolution $81922^{13}=$ 8192, where lower resolution data are down-sampled from.

### v3

@@ 3.2 Compensation Methods of Low-Frequency Tendency
In order to address the issues caused by slow convergence at high frequencies and speed up the training of FNO, several methods have been proposed. Jagtap and Ameya D [27] used activation functions $\sigma(\mu ax)$ instead of $\sigma(\mu x)$. Here, $\mu$ is a fixed scaling factor, and $a$ is a trainable parameter for all neurons. Cai and Wei [28] introduced PhaseDNN, which converts the high-frequency components of data to the low-frequency spectrum for learning, and then converts the learned frequency components back to the original high-frequency spectrum. Subsequently, the multiscale DNN method (MscaleDNN) was proposed [22, 29], which can be understood as Fourier expansion only in the radial direction. The frequency space conversion can be achieved through scaling, which is equivalent to inverse scaling in spatial space. In this article, we utilize the conversion of coarse and fine grids and MscaleDNN to address the slow convergence issue of high-frequency errors in FNO. Next, let us review the ideas and structure of MscaleDNN in regular DNNs.
@@
A common MscaleDNN structure is shown in Figure 6. We use neural networks successively in physical space to learn the stretched sub-functions $f_{i}^{(scale)}(\alpha_{i}x)$. Furthermore, it is mentioned in [Liu_2020] that when constructing MscaleDNN, activation functions with compact support should be selected. In MgFNO, we select the following function $\phi(x)$ as our activation function.
@@
Figure 6: : A common MscaleDNN structure
On the other hand, looking back at the MG method we mentioned in Section 2, it is worth noting that iterative methods (2.1) tend to eliminate high-frequency error components rather than low-frequency ones. Therefore, we may employ a straightforward approach to remove these low-frequency error components. As it is shown in figure 7, for MG method on a coarser mesh, the low frequency components transform into high frequency ones. Because if there is an error on one side of the mesh and a different error on the other side of the mesh, once we coarsen the cells and then bring them next to each other, the variation between those two positions is now a high spatial frequency. Then we can remove high-frequency components with a few iterations of iterative methods. Due to the opposite convergence tendency of DNNs and iterative methods in the frequency domain, we have decided to apply a multi-grid architecture similar to the MG method to FNO, but this time we need to transfer the coarse grid to the fine grid. We hope that through the design of this multi-grid architecture, FNO can overcome the high-frequency issues and accelerate the model training process.
@@ 4.2 Architecture of MgFNO
Due to the FNO adhering to the F-Principle, we suggest that mitigating the high-frequency errors of FNO could further enhance its overall performance. To accomplish this, we propose MgFNO, a multigrid framework that trains multiple FNOs sequentially. There are various multigrid algorithms available, such as the V-cycle and W-cycle. For the V-cycle variant, the model employs a three-level hierarchy where: the coarse-level network,the intermediate network and the fine-level network.Training follows a strict bottom-up progression with parameter transfer between levels. The W-cycle modification introduces repeated coarsening-refinement steps at each level, particularly enhancing low-frequency error correction - while increasing computational cost by ∼50% compared to V-cycle, it demonstrates superior convergence for non-elliptic problems like the Navier-Stokes equations. Both variants maintain the FNO’s resolution invariance while accelerating convergence through this hierarchical frequency decomposition. For the convenience of model construction, we employ a three-layer V-cycle multigrid architecture to build the model, and the specific algorithm of MgFNO is depicted in Algorithm 2.
@@ 5 Numerical Implementation and Results
Table 2 presents a comparison between the proposed MgFNO and the standard FNO in terms of parameter count and the required training epochs for MgFNO to achieve the final error level of standard FNO. The results demonstrate that while MgFNO’s parameter count is three times that of standard FNO due to the construction of three baseline networks, it only requires 25% of the training epochs needed by the conventional approach owing to its effective mitigation of the frequency principle.
@@
Table 2: The number of parameters for different network structures and the epochs required for each model.
Error
FNO MgFNO
Parameters/Epochs Parameters/Epochs
Burgers $1.60\%\pm 0.5\%$ 582849/600 1748547/50+50+50
Darcy $0.9\%\pm 0.5\%$ 2368001/600 7104003/50+50+50
Navier-Stokes $1.28\%\pm 0.5\%$ 928661/600 2785983/60+60+60
@@
Data Set and Result. Firstly, we set the viscosity coefficient $\nu=0.1$. Then, for initial condition $u_{0}(x)$, we set $u_{0}\sim\mu$ where $\mu=\mathcal{N}(0,625(-\Delta+25I)^{-2})$. Finally, we solve the equation using a split step method where the heat equation part is solved exactly in Fourier space then the non-linear part is advanced, again in Fourier space, using a very fine forward Euler method. We solve on a spatial mesh with resolution $81922^{13}=$ 8192, where lower resolution data are downsampled from.
@@
Frequency analysis.
(a) FNO in frequency domain.
(b) MgFNO in frequency domain.
Figure 12: :Comparative analysis of convergence behavior between MgFNO and FNO during training in frequency domain.
In this section, we present experimental evidence demonstrating how the proposed MgFNO effectively overcomes the inherent limitation of standard FNO in converging high-frequency errors through its multigrid architecture. Figure 12 illustrates the comparative frequency-domain convergence between FNO and MgFNO during training, with the MgFNO transitioning to its second and third grid levels at iteration steps 4000 and 8000, respectively. The results reveal that while conventional FNO sequentially converges low-frequency components before addressing high-frequency ones - a characteristic manifestation of the frequency principle - it consistently fails to adequately learn high-frequency components, resulting in significant residual errors. In contrast, our MgFNO strategically performs grid transitions when low-frequency learning stagnates: the high-frequency components that prove challenging to learn on finer grids are effectively converted into more tractable low-frequency representations on coarser grids. This adaptive multiscale approach enables MgFNO to achieve superior approximation accuracy of the target operator, with experimental measurements showing a 42% reduction in high-frequency error compared to standard FNO under identical training budgets. The architectural innovation fundamentally rebalances the frequency learning priority while maintaining the resolution-invariant property of neural operators.
@@ 6 Conclusions
Acknowledgements. The authors are very much indebted to the referees for their constructive comments and valuable suggestions.
CONFLICT OF INTEREST STATEMENT
This study does not have any conflicts to disclose.
DATA AVAILABILITY
Code and data used are available on https://github.com/guozihao-hub/MgFNO/tree/master.